# Boundary phase transitions of two-dimensional quantum critical XXZ model

### v1

@@ 4 Irrelevance of weak velocity difference
4.1 The dangling spin chain case
@@
4.2 The dangling two-leg ladder case
@@
$\begin{split}\int\mathrm{d}z\mathrm{d}\bar{z}\,\sum_{i=\sigma,\tau}\bigg{(}&-\frac{\pi^{2}}{16}(3-2\epsilon_{xx})(g_{xx}^{i})^{2}\delta v\mathrm{d}l(\partial_{z}\theta_{\mathrm{S}}\partial_{z}\theta_{\mathrm{S}}+\partial_{\bar{z}}\theta_{\mathrm{S}}\partial_{\bar{z}}\theta_{\mathrm{S}})\\ &-\frac{\pi^{2}}{8}(3-2\epsilon_{z})(g^{i}_{z})^{2}\delta v\mathrm{d}l(\partial_{z}\phi_{\mathrm{S}}\partial_{z}\phi_{\mathrm{S}}+\partial_{\bar{z}}\phi_{\mathrm{S}}\partial_{\bar{z}}\phi_{\mathrm{S}})\bigg{)},\end{split}$ (87)
@@
Table 1: Summary of possible boundary phases of the 2D quantum critical AF XXZ model weakly coupled with either a dangling spin-$1/2$ XXZ chain or a two-leg ladder on the boundary.
easy-plane AF spin rotation O(2) ordinary extraordinary
VBS lattice translation ordinary ordinary
easy-plane AF spin rotation O(2) ordinary extraordinary
singlet none ordinary ordinary
5 Summary
We studied the boundary critical behavior of a 2D quantum critical AF XXZ model coupled either to a dangling spin chain or a two-leg ladder on its boundary. In each case, we start from the bosonized effective field theory of the spin chain or the ladder, and examine the stability of phases in the presence of the coupling to critical bulk AF modes with the perturbative RG analysis. We find a rich boundary phase diagram in each case. The main results are summarized in Table 1.
6 Acknowledgements
@@ B.2 The dangling two-leg ladder case
They reduce to Eqs. (65)–(73) and lead to the three phases discussed in Sec. 3 if the $\gamma_{\phi}$ and $\gamma_{\theta}$ terms are neglected. In these phases, the extra terms in the RG equations produced by $\gamma_{\phi}$ and $\gamma_{\theta}$ turn out to be subleading and do not change the asymptotic RG flow of the other coupling constants. Therefore, these boundary phases are stable against the mixing of boson modes in the two sectors.

### v3

@@ IV Irrelevance of weak velocity difference
IV.1 The dangling spin chain case
@@
IV.2 The dangling two-leg ladder case
@@
$\int\mathrm{d}z\mathrm{d}\bar{z}\,\sum_{i=\sigma,\tau}\bigg{(}-\frac{\pi^{2}}{16}(3-2\epsilon_{xx})(g_{xx}^{i})^{2}\delta v\mathrm{d}l(\partial_{z}\theta_{\mathrm{S}}\partial_{z}\theta_{\mathrm{S}}+\partial_{\bar{z}}\theta_{\mathrm{S}}\partial_{\bar{z}}\theta_{\mathrm{S}})-\frac{\pi^{2}}{8}(3-2\epsilon_{z})(g^{i}_{z})^{2}\delta v\mathrm{d}l(\partial_{z}\phi_{\mathrm{S}}\partial_{z}\phi_{\mathrm{S}}+\partial_{\bar{z}}\phi_{\mathrm{S}}\partial_{\bar{z}}\phi_{\mathrm{S}})\bigg{)},$ (90)
@@
V Conclusion and discussions
We have studied the boundary critical behavior of a 2D quantum critical AF XXZ model coupled either to a dangling spin chain or a two-leg ladder on its boundary. In each case, we start from the bosonized effective field theory of the spin chain or the ladder, and examine the stability of phases in the presence of the coupling to critical bulk AF modes with the perturbative RG analysis. The main results are summarized in Tables 1 and 2.
The RG analysis reveals a rich boundary phase diagram of the (2+1)D quantum XXZ model at the bulk QCP. In the easy-axis case, the boundary can be either in the ordinary universality class or in the extraordinary class with an Ising AF order on the boundary, which is the same as the 3D classical Ising model. However, the easy-plane and the isotropic cases are quite different from the 3D classical counterparts. The dangling spin chain coupled to the critical bulk state either breaks the lattice translation symmetry by forming a VBS order or breaks the spin rotation symmetry by forming an AF order, which generalizes the conclusion of the isotropic case in Ref. [27]. The dangling two-leg ladder coupled to the critical bulk state can either form a symmetric boundary with ordinary critical behavior or show an AF order with extraordinary critical behavior. When the two-leg ladder is gapped with only short-range correlation, a weak coupling to the bulk is irrelevant and does not change the boundary critical behavior. On the other hand, if the ladder is critical with a quasi-long-range spin correlation, the coupling to the bulk is relevant and results into a long-range AF order and extraordinary critical behavior on the boundary. The boundary AF order and a direct transition to the VBS order in the dangling spin chain case and to the symmetric boundary state in the dangling two-leg ladder case have been observed in numerical simulations [28, 30], which is captured by the RG analysis in Ref. [27] and this work.
The boundary critical behavior is closely related to the boundary conformal field theory (BCFT). Our RG analysis suggests that coupling a low-dimensional quantum system to the boundary of a quantum critical state can engineer new universality classes of boundary critical behavior, thereby it provides a novel approach to study the BCFT, which is complementary to the bootstrap and the holographic method [45, 46, 47, 48].
Our work provides a few concrete examples of RG flow on the boundary of a QCP. While the boundary RG flow of the 2D conformal field theory is dictated by the $g$-theorem, which states that the Affleck-Ludwig boundary entropy monotonically decreases with the RG flow [49, 50], its counterpart in higher dimensions is by far lacking. This will be investigated in our future numerical and analytic works on the boundary critical behavior.
VI Acknowledgements
@@
VII Data availability
No data were created or analyzed in this study.
@@ B.2 The dangling two-leg ladder case
They reduce to Eqs. (68)–(76) and lead to the three phases discussed in Sec. III if the $\gamma_{\phi}$ and $\gamma_{\theta}$ terms are neglected. In these phases, the extra terms in the RG equations produced by $\gamma_{\phi}$ and $\gamma_{\theta}$ turn out to be subleading and do not change the asymptotic RG flow of the other coupling constants. Therefore, these boundary phases are stable against the mixing of boson modes in the two sectors.